# Strong semistability of Higgs bundles over curves

### v4

@@ 4. Semistable but not strongly semistable Higgs bundles
We first provide a construction of semistable but not strongly semistable Higgs bundle for each big rank, thanks to the following result of X. Sun.
@@
$0=\mathcal{V}^{n(p-1)+1}\subset\mathcal{V}^{n(p-1)}\subset\dots\subset\mathcal{V}^{0}=\mathcal{V},$
@@
$\mathcal{V}^{\ell+1}=\ker\{\mathcal{V}^{\ell}\stackrel{{\scriptstyle\nabla}}{{\longrightarrow}}\mathcal{V}\otimes K_{C}\to\mathcal{V}/\mathcal{V}^{\ell}\otimes K_{C}\}.$
@@ 1. $(1)$
    $\mathcal{V}^{0}/\mathcal{V}^{1}\cong\mathcal{E},\nabla(\mathcal{V}^{\ell+1})\subseteq\mathcal{V}^{\ell}\otimes K_{C}$ for $\ell\geq 1$.
@@ 2. $(2)$
    $\mathcal{V}^{\ell}/\mathcal{V}^{\ell+1}\stackrel{{\scriptstyle\nabla}}{{\longrightarrow}}(\mathcal{V}^{\ell-1}/\mathcal{V}^{\ell})\otimes K_{C}$ is an isomorphism for $1\leq\ell\leq p-1$.
@@ 3. $(3)$
From now on till end, we assume the genus $g(C)\geq 2$. Let $\mathcal{L}$ be a line bundle on $C$ and $\mathcal{E}_{1}=F_{*}\mathcal{L}$. By Theorem 4.1, $\mathcal{E}_{1}$ is a stable rank $p$ bundle of degree $\deg\mathcal{L}+(p-1)(g-1)$. Moreover, by Lemma 4.1 the canonical filtration of $F^{*}\mathcal{E}_{1}$ is a gr-semistable filtration, and it is unique since the associated graded Higgs bundle with respect to the canonical filtration is
@@
where $\theta_{i}\colon K_{C}^{i}\to K_{C}^{i-1}\otimes K_{C}$ is an isomorphism for each $i=1,\dots,p-1$.
If we choose $\mathcal{L}=F^{*}\mathcal{L}_{2}$ for some line bundle $\mathcal{L}_{2}$, then $\mathcal{L}\otimes K_{C}^{p}=F^{*}(\mathcal{L}_{2}\otimes K_{C})$ can be equipped with the canonical connection. Now we define $\mathcal{V}=\left(\mathcal{L}\otimes K_{C}^{p}\right)\oplus F^{*}\mathcal{E}_{1}$ together with a flat connection $\nabla$ given by
$\begin{pmatrix}\nabla_{\operatorname{can}}&\theta_{p}\\ 0&\nabla_{\operatorname{can}}\end{pmatrix}$,
where $\theta_{p}\colon\mathcal{L}\otimes K_{C}^{p}\stackrel{{\scriptstyle\cong}}{{\to}}\mathcal{L}\otimes K_{C}^{p-1}\otimes K_{C}\subseteq F^{*}\mathcal{E}_{1}\otimes K_{C}$ is the natural morphism.
@@
It is clear that the exponent of nilpotency of $p$-curvature of $(\mathcal{V},\nabla)$ is $\leq 1$, since $0\subset F^{*}\mathcal{E}_{1}\subset\mathcal{V}$ is a filtration of $\mathcal{V}$ such that the $p$-curvature of each graded piece is zero. Thus $(\mathcal{E},\theta)=C(\mathcal{V},\nabla)$ is well-defined, and $(\mathcal{E},\theta)$ is a stable Higgs bundle, since $(\mathcal{V},\nabla)$ is $\nabla$-stable.
@@
The stable rank $p+1$ Higgs bundle $(\mathcal{E},\theta)$ constructed as above is not strongly semistable.
@@
Note that the gr-semistable filtration on $(\mathcal{V},\nabla)=C^{-1}(\mathcal{E},\theta)$ is given by
$0\subset\mathcal{L}\otimes K_{C}^{p}\subset\mathcal{L}\otimes K_{C}^{p}\oplus\mathcal{N}^{\bullet},$
where $\mathcal{N}^{\bullet}$ is the canonical filtration of $F^{*}\mathcal{E}_{1}$. Moreover, the gr-semistable filtration of $(\mathcal{V},\nabla)$ is unique since the associated graded Higgs bundle is
$(K_{C}^{p}\oplus K_{C}^{p-1}\oplus K_{C}^{p-2}\oplus\dots\oplus\mathcal{O}_{C},\bigoplus_{i=1}^{p}\theta_{i})\otimes(\mathcal{L},0),$
which is a stable Higgs bundle with the exponent of nilpotency =p=p. By Lemma 2.3, there is no other gr-semistable filtration on $(\mathcal{V},\nabla)$ such that the exponent of nilpotency of the associated graded Higgs bundle is $\leq p-1$. This completes the proof. ∎
@@
Let $C$ be a curve of genus $g\geq 2$. There exists a semistable but not strongly semistable Higgs bundle on $C$ of any big rank.
@@
By Proposition 4.1, there exists a semistable Higgs bundle $(\mathcal{E},\theta)$ which is not strongly semistable of rank $p+1$. By construction of $(\mathcal{E},\theta)$ we have the slope of $\mathcal{E}$ is an integer. Let $\mathcal{M}$ be a line bundle such that $\deg\mathcal{M}=\mu(\mathcal{E})$. Then $(\widetilde{\mathcal{E}},\widetilde{\theta})=(\mathcal{E},\theta)\oplus(\mathcal{M},0)$ is a semistable Higgs bundle. Since the inverse Cartier transform preserves direct sum,
@@
$0\subset\mathcal{M}^{p}\subset\mathcal{M}^{p}\oplus\mathcal{L}\otimes K_{C}^{p}\subset\mathcal{M}^{p}\oplus\mathcal{L}\otimes K_{C}^{p}\oplus\mathcal{N}^{\bullet},$
@@
Theorem 4.2 ([LSYZ19, Corollary 2.2]).
@@
$(\mathrm{Gr}_{\text{HN}}\circ C^{-1})(\mathcal{E}_{\operatorname{unif}},\theta_{\operatorname{unif}})\cong(\mathcal{E}_{\operatorname{unif}},\theta_{\operatorname{unif}})\otimes(\mathcal{L},0),$
where HN denotes for the Harder-Narasimhan filtration, and $\mathcal{L}$ is a torsion line bundle of order two.
Put $(\mathcal{V}_{\operatorname{unif}},\nabla_{\operatorname{unif}})=C^{-1}\circ\mathrm{Gr}_{\text{HN}}\circ C^{-1}(\mathcal{E}_{\operatorname{unif}},\theta_{\operatorname{unif}})$. Then by Theorem 4.2, it is a uniformizing flat bundle with nilpotent $p$-curvature. For the case $p=2$, the above theorem does not work. However, there is a direct way to construct a uniformizing flat bundle when $g(C)$ is odd: Since both $K_{C}\cong F^{*}K_{C}^{\frac{1}{2}}$ and $\mathcal{O}_{C}$ can be equipped with the canonical connection, we define $\mathcal{V}=K_{C}\oplus\mathcal{O}_{C}$ and equip it with a flat connection $\nabla$ defined by
$\begin{pmatrix}\nabla_{\operatorname{can}}&\theta\\ 0&\nabla_{\operatorname{can}}\end{pmatrix}$, (4.1)
@@
Clearly, $(\mathcal{E}_{1},\theta_{1})$ is a (nilpotent) stable Higgs bundle. Next, let $\mathcal{E}_{2}=F_{*}K_{C}^{\frac{1-p}{2}}$. By Theorem 4.1, $\mathcal{E}_{2}$ is a stable bundle of rank $p$ and degree zero. Moreover, by Lemma 4.1 the unique gr-semistable filtration on $F^{*}\mathcal{E}_{2}$ is given by its canonical filtration, and the associated graded Higgs bundle is
@@
where $\theta_{i}\colon K_{C}^{i}\to K_{C}^{i-1}\otimes K_{C}$ is an isomorphism for each $i=\frac{3-p}{2},\dots,\frac{p-1}{2}$. Third, we show that $(\mathcal{E}_{1},\theta_{1})\otimes(\mathcal{E}_{2},0)$ is a stable Higgs bundle. For $p=2$, as $(\mathcal{O}_{C},0)$ is the only Higgs subbundle of $(\mathcal{E}_{1},\theta_{1})$, all Higgs subbundles of the tensor product are of form $(\mathcal{O}_{C}\otimes\mathcal{F},0)$, where $\mathcal{F}$ is a subbundle of $\mathcal{E}_{2}$. The stability of the tensor product follows immediately. For $p$ odd, this is more than obvious, since we have explicitly
@@
which is clearly a stable Higgs bundle. Fourth, there is the following diagram of the Higgs-de Rham flow operator acting upon $(\mathcal{E}_{1},\theta_{1})\otimes(\mathcal{E}_{2},0)$:
$C^{-1}(\mathcal{E}_{1},\theta_{1})\otimes(F^{*}F_{*}K_{C}^{\frac{1-p}{2}},\nabla_{\operatorname{can}})$$(\mathcal{E}_{1},\theta_{1})\otimes(\mathcal{E}_{2},0)$$(\mathcal{E}^{\prime},\theta^{\prime})$$\scriptstyle{\mathcal{N}^{\bullet}}$$\scriptstyle{C^{-1}}$
@@
By Lemma 4.2 below, $(\mathcal{E}^{\prime},\theta^{\prime})$ is polystable whose exponent of nilpotency is exactly $p$. Hence $\mathcal{N}^{\bullet}$ is a gr-polystable filtration. It follows from Lemma 2.3 that there is no gr-semistable filtration on $C^{-1}(\mathcal{E}_{1},\theta_{1})\otimes(F^{*}F_{*}K_{C}^{\frac{1-p}{2}},\nabla_{\operatorname{can}})$ such that its associated graded Higgs bundle has exponent of nilpotency $\leq p-1$. This shows that $(\mathcal{E}_{1},\theta_{1})\otimes(\mathcal{E}_{2},0)$ is not strongly semistable. Fifth, by the strongly semistability theorem, both $(\mathcal{E}_{1},\theta_{1})$ and $(\mathcal{E}_{2},0)$ are strongly semistable. Summarizing the above altogether, we have constructed an example of strongly semistable Higgs bundles whose tensor product is semistable but not strongly semistable, as claimed.
@@
Firstly we prove that for each $m\in\operatorname{\mathbb{Z}}_{>0}$, the $m$-th symmetric product of the uniformizing Higgs bundle
$\operatorname{Sym}^{m}(\mathcal{E}_{\operatorname{unif}},\theta_{\operatorname{unif}})=(K_{C}^{\frac{m}{2}}\oplus\dots\oplus K_{C}^{-\frac{m}{2}},\bigoplus_{i=1}^{m}\theta_{i})$
is stable. By restricting to the generic point, this reduces to a linear algebra problem, that is, a $k(C)$-vector space $V$ of dimension $m+1$ together with a basis $\{e_{1},\dots,e_{m+1}\}$ such that
$\theta(e_{1},\dots,e_{m+1})=(e_{1},\dots,e_{m+1})J_{m+1}(0),$
As a consequence, the rank $r$ Higgs subbundle of $\operatorname{Sym}^{m}(\mathcal{E}_{\operatorname{unif}},\theta_{\operatorname{unif}})$ must be the direct sum of its last $r$ factors, which is of negative degree, and this shows $\operatorname{Sym}^{m}(\mathcal{E}_{\operatorname{unif}},\theta_{\operatorname{unif}})$ is a stable Higgs bundle.

### v5

@@ 4. Semistable but not strongly semistable Higgs bundles
We first provide a construction of semistable but not strongly semistable Higgs bundle of any big rank, thanks to the following results of X. Sun.
@@
Theorem 4.2 ([Sun10, Corollary 2.4]).
Let $C$ be a smooth projective curve of genus $g\geq 2$ and $\mathcal{L}$ be a line bundle on $C$. Then, for all proper subbundles $\mathcal{E}\subseteq F_{*}\mathcal{L}$, we have
$\mu(\mathcal{E})-\mu(F_{*}\mathcal{L})\leq-\frac{p-\mathrm{rk}(\mathcal{E})}{p}(g-1).$
@@
$\mathcal{V}^{p}\subset\mathcal{V}^{p-1}\subset\dots\subset\mathcal{V}^{0}=\mathcal{V},$
@@
$\mathcal{V}^{\ell+1}=\ker\{\mathcal{V}^{\ell}\stackrel{{\scriptstyle\nabla_{\operatorname{can}}}}{{\longrightarrow}}\mathcal{V}\otimes K_{C}\to\mathcal{V}/\mathcal{V}^{\ell}\otimes K_{C}\}.$
@@
    $\mathcal{V}^{0}/\mathcal{V}^{1}\cong\mathcal{E},\nabla_{\operatorname{can}}(\mathcal{V}^{\ell+1})\subseteq\mathcal{V}^{\ell}\otimes K_{C}$ for $\ell\geq 1$.
@@
    $\mathcal{V}^{\ell}/\mathcal{V}^{\ell+1}\stackrel{{\scriptstyle\nabla_{\operatorname{can}}}}{{\longrightarrow}}(\mathcal{V}^{\ell-1}/\mathcal{V}^{\ell})\otimes K_{C}$ is an isomorphism for $1\leq\ell\leq p-1$ and $\mathcal{V}^{p}=0$.
@@
From now on till end, we assume $g\geq 2$. Let $\mathcal{L}$ be a line bundle on $C$ and $\mathcal{E}^{\prime}=F_{*}\mathcal{L}$. By Theorem 4.1 and Riemann-Roch theorem, $\mathcal{E}^{\prime}$ is a stable rank $p$ bundle of degree $\deg\mathcal{L}+(p-1)(g-1)$. Moreover, by Lemma 4.1 the canonical filtration of $F^{*}\mathcal{E}^{\prime}$ is a Griffiths transverse saturated filtration, and by Lemma 2.2 it is the unique gr-semistable filtration since the associated graded Higgs bundle with respect to the canonical filtration is
@@
which is a stable Higgs bundle by Lemma 4.2 below, where $\theta_{i}\colon K_{C}^{i}\to K_{C}^{i-1}\otimes K_{C}$ is an isomorphism for each $i=1,\dots,p-1$.
Lemma 4.2.
For each $m\in\operatorname{\mathbb{Z}}_{>0}$, the Higgs bundle
$(\mathcal{F}^{m},\theta^{m})=(K_{C}^{m}\oplus K_{C}^{m-1}\oplus\dots\oplus K_{C}\oplus\mathcal{O}_{C},\bigoplus_{i=1}^{m}\theta_{i})$
is stable, where $\theta_{i}$ is an isomorphism for each $i=1,\dots,m$.
Proof.
By restricting to the generic point, this reduces to a linear algebra problem, that is, a $k(C)$-vector space $V$ of dimension $m+1$ together with a basis $\{e_{1},\dots,e_{m+1}\}$ such that
$\theta^{m}(e_{1},\dots,e_{m+1})=(e_{1},\dots,e_{m+1})J_{m+1}(0),$
As a consequence, the rank $r$ Higgs subbundle of $(\mathcal{F}^{m},\theta^{m})$ must be the direct sum of its last $r$ factors, whose slope is strictly less than the slope of $\mathcal{F}^{m}$, and this shows $(\mathcal{F}^{m},\theta^{m})$ is a stable Higgs bundle. ∎
For each $0<\ell\leq 2(g-1)$, we choose and then fix a degree $\ell$ line bundle $\mathcal{A}_{\ell}$ such that $H^{0}(C,K_{C}\otimes\mathcal{A}_{\ell}^{-1})\neq 0$. Fix $x_{0}\in C(k)$, we take $\mathcal{L}$ in the argument above to be some line bundle $\mathcal{L}_{\ell}$ such that
$\mathcal{A}_{\ell}\otimes\mathcal{L}_{\ell}\otimes K_{C}^{p-1}\cong\mathcal{O}_{C}(mpx_{0})$
for some $m\in\operatorname{\mathbb{Z}}$. Then $\mathcal{A}_{\ell}\otimes\mathcal{L}_{\ell}\otimes K_{C}^{p-1}$ can be equipped with the canonical connection. Pick a non-zero element in $H^{0}(C,K_{C}\otimes\mathcal{A}_{\ell}^{-1})$, which defines a non-zero morphism
$\theta\colon\mathcal{A}_{\ell}\otimes\mathcal{L}_{\ell}\otimes K_{C}^{p-1}\to\mathcal{L}_{\ell}\otimes K_{C}^{p-1}\otimes K_{C}\subseteq F^{*}\mathcal{E}^{\prime}\otimes K_{C}.$
Set
$\mathcal{V}_{\ell}=\mathcal{A}_{\ell}\otimes\mathcal{L}_{\ell}\otimes K_{C}^{p-1}\oplus F^{*}\mathcal{E}^{\prime},$
which is equipped with a connection $\nabla_{\ell}$ given by
$\begin{pmatrix}\nabla_{\operatorname{can}}&\theta\\ 0&\nabla_{\operatorname{can}}\end{pmatrix}$,
where $\theta$ is defined as above.
Lemma 4.3.
$(\mathcal{V}_{\ell},\nabla_{\ell})$ is $\nabla$-stable.
Proof.
By Riemann-Roch theorem, we know that
$\deg(F^{*}\mathcal{E}^{\prime})=p\left((p-1)(g-1)+\deg(\mathcal{L}_{\ell})\right),$
and hence
$\deg(\mathcal{V}_{\ell})=(p+2)(p-1)(g-1)+(p+1)\deg(\mathcal{L}_{\ell})+\ell.$
Thus
$\mu(F^{*}\mathcal{E}^{\prime})=(p-1)(g-1)+\deg(\mathcal{L}_{\ell})<\frac{p+2}{p+1}(p-1)(g-1)+\frac{\ell}{p+1}+\deg(\mathcal{L}_{\ell})=\mu(\mathcal{V}_{\ell}).$
Now suppose $\mathcal{M}$ is a $\nabla$-invariant proper subbundle of $(\mathcal{V}_{\ell},\nabla_{\ell})$. Consider the composition
$\mathcal{M}\hookrightarrow\mathcal{V}_{\ell}\stackrel{{\scriptstyle\pi}}{{\longrightarrow}}\mathcal{A}_{\ell}\otimes\mathcal{L}_{\ell}\otimes K_{C}^{p-1},$
where $\pi$ is the natural projection. There are two cases:
  • $(1)$
    $\pi(\mathcal{M})=0$: In this case we have $\mathcal{M}\subseteq F^{*}\mathcal{E}^{\prime}$. By Cartier descent, $(F^{*}\mathcal{E}^{\prime},\nabla_{\operatorname{can}})$ is a $\nabla$-stable bundle as $\mathcal{E}^{\prime}$ is a stable bundle. Hence
    $\mu(\mathcal{M})<\mu(F^{*}\mathcal{E}^{\prime})<\mu(\mathcal{V}_{\ell}).$
  • $(2)$
    $\pi(\mathcal{M})=\mathcal{A}_{\ell}\otimes\mathcal{L}_{\ell}\otimes K_{C}^{p-1}$: In this case we have the following exact sequence:
    $0\to\mathcal{N}\to\mathcal{M}\to\mathcal{A}_{\ell}\otimes\mathcal{L}_{\ell}\otimes K_{C}^{p-1}\to 0,$
    where $\mathcal{N}=\mathcal{M}\cap F^{*}\mathcal{E}^{\prime}$ is a $\nabla$-invariant subbundle of $F^{*}\mathcal{E}^{\prime}$. Claim that $\mathcal{N}\neq 0$. Otherwise, we have $\mathcal{M}\cong\mathcal{A}_{\ell}\otimes\mathcal{L}_{\ell}\otimes K_{C}^{p-1}$. As $\deg(\mathcal{A}_{\ell})=\ell>0$, $\mathcal{A}_{\ell}\otimes\mathcal{L}_{\ell}\otimes K_{C}^{p-1}$ is actually the maximal destabilizer of $\mathcal{V}_{\ell}$. Therefore, $\mathcal{M}=\mathcal{A}_{\ell}\otimes\mathcal{L}_{\ell}\otimes K_{C}^{p-1}$. However, it is not $\nabla$-invariant by construction. Contradiction. Denote the rank of $\mathcal{N}$ by $n$. Now we know $0<n<p$. Again by Cartier descent, there is a unique rank $n$ subbundle $\widetilde{\mathcal{N}}$ of $\mathcal{E}^{\prime}$ such that $\mathcal{N}=F^{*}\widetilde{\mathcal{N}}$. By Theorem 4.2, we know that
    $\frac{\deg(\mathcal{N})}{np}=\frac{\deg(\widetilde{\mathcal{N}})}{n}\leq\frac{(g-1)}{p}\left(n-1\right)+\frac{\deg(\mathcal{L}_{\ell})}{p}.$
    Thus
    $\mu(\mathcal{M})$ $=\frac{\deg(\mathcal{N})+2(p-1)(g-1)+\deg(\mathcal{L}_{\ell})+\ell}{n+1}$
    $\leq\frac{\left(n^{2}-n+2p-2\right)(g-1)+\ell}{n+1}+\deg(\mathcal{L}_{\ell}).$
    Now we show if $0<n<p$ and $0<\ell\leq 2(g-1)$, then $\mu(\mathcal{M})\leq\mu(\mathcal{V}_{\ell})$. Indeed, it suffices to show
    $\frac{\left(n^{2}-n+2p-2\right)(g-1)+\ell}{n+1}\leq\frac{p+2}{p+1}(p-1)(g-1)+\frac{\ell}{p+1},$
    which is reduced to show
    $\frac{(p-n)}{(p+1)(n+1)}(n(p+1)(g-1)-(p-1)(g-1)-\ell)\geq 0.$
    But this is clear. Furthermore, when $\ell<2(g-1)$, the above inequality is strict. It implies $(\mathcal{V}_{\ell},\nabla_{\ell})$ is $\nabla$-stable if $\ell<2(g-1)$. Finally we show that $(\mathcal{V}_{\ell},\nabla_{\ell})$ is also $\nabla$-stable when $\ell=2(g-1)$. Suppose $\ell=2(g-1)$. (In this case, $\mathcal{A}_{\ell}=K_{C}$.) The inequality above becomes an equality only when $\mathrm{rk}(\mathcal{N})=1$. By above, $\deg(\mathcal{N})\leq\deg(\mathcal{L}_{\ell})$. Without lose of generality we may simply take $\mathcal{L}_{\ell}=\mathcal{O}_{C}$ in the case $\ell=2(g-1)$. If $\deg(\mathcal{N})<0$, then
    $\mu(\mathcal{M})<p(g-1)=\mu(\mathcal{V}_{2(g-1)}).$
    So we may assume $\deg(\mathcal{N})=0$. Notice that $\mathcal{M}$ cannot contain $K_{C}^{p}$: If it were the case, since
    $K_{C}^{p}\stackrel{{\scriptstyle\nabla|_{K_{C}^{p}}}}{{\longrightarrow}}(K_{C}^{p}\oplus K_{C}^{p-1})\otimes K_{C}\twoheadrightarrow K_{C}^{p-1}\otimes K_{C}=K_{C}^{p}$
    is an isomorphism, $\mathcal{M}$ would contain $K_{C}^{p}\oplus K_{C}^{p-1}$. Then, since
    $K_{C}^{p-1}\stackrel{{\scriptstyle\nabla|_{K_{C}^{p-1}}}}{{\longrightarrow}}\mathcal{N}^{p-2}\otimes K_{C}\twoheadrightarrow K_{C}^{p-2}\otimes K_{C}=K_{C}^{p-1}$
    is again an isomorphism by Lemma 4.1, where $\mathcal{N}^{\bullet}$ is the canonical filtration on $F^{*}\mathcal{E}^{\prime}$. This shows $\mathcal{M}$ would contain $K_{C}^{p}\oplus\mathcal{N}^{p-2}$, which is impossible. Thus, if we define a filtration on $(\mathcal{V}_{2(g-1)},\nabla_{2(g-1)})$ by
    $0\subset K_{C}^{p}\subset K_{C}^{p}\oplus\mathcal{N}^{\bullet},$
    it follows that
    $\mathrm{Gr}(\mathcal{M},\nabla_{2(g-1)}|_{\mathcal{M}})\subset\mathrm{Gr}(F^{*}\mathcal{E}^{\prime},\nabla_{\operatorname{can}})$
    where $\mathrm{Gr}$ denotes the graded bundle associated to the above filtration. Since $\mu(\mathcal{M})=p(g-1)>(p-1)(g-1)=\mu(F^{*}\mathcal{E}^{\prime})$, we derive a contradiction to the stability of $\mathrm{Gr}(F^{*}\mathcal{E}^{\prime},\nabla_{\operatorname{can}})$ as asserted in Lemma 4.2. It implies that such a $\nabla$-invariant rank $2$ subbundle $\mathcal{M}$ which is an extension of $K_{C}^{p}$ by a degree zero line bundle $\mathcal{N}\subset F^{*}\mathcal{E}^{\prime}$ does not exist. So $(\mathcal{V}_{2(g-1)},\nabla_{2(g-1)})$ is also $\nabla$-stable. This completes the proof.
@@
It is clear that the exponent of nilpotency of $p$-curvature of $(\mathcal{V}_{\ell},\nabla_{\ell})$ is $\leq 1$, since $0\subset F^{*}\mathcal{E}^{\prime}\subset\mathcal{V}_{\ell}$ is a filtration of $\mathcal{V}_{\ell}$ such that the $p$-curvature of each graded piece is zero. Thus $(\mathcal{E}_{\ell},\theta_{\ell})=C(\mathcal{V}_{\ell},\nabla_{\ell})$ is well-defined, and it is a stable Higgs bundle as followed from Lemma 4.3.
@@
For integer $\ell$ with $0<\ell\leq 2(g-1)$, the nilpotent stable rank $p+1$ Higgs bundle $(\mathcal{E}_{\ell},\theta_{\ell})$ of exponent $\leq 1$ constructed as above is not strongly semistable.
@@
Note that a Griffiths transverse saturated filtration on $(\mathcal{V}_{\ell},\nabla_{\ell})=C^{-1}(\mathcal{E}_{\ell},\theta_{\ell})$ is given by
$0\subset\mathcal{A}_{\ell}\otimes\mathcal{L}_{\ell}\otimes K_{C}^{p-1}\subset\mathcal{A}_{\ell}\otimes\mathcal{L}_{\ell}\otimes K_{C}^{p-1}\oplus\mathcal{N}^{\bullet},$
where $\mathcal{N}^{\bullet}$ is the canonical filtration of $F^{*}\mathcal{E}^{\prime}$. Moreover, this filtration of $(\mathcal{V}_{\ell},\nabla_{\ell})$ is the unique gr-semistable one since the associated graded Higgs bundle is
$(\mathcal{A}_{\ell}\otimes K_{C}^{p-1}\oplus K_{C}^{p-1}\oplus K_{C}^{p-2}\oplus\dots\oplus\mathcal{O}_{C},\theta\oplus\bigoplus_{i=1}^{p-1}\theta_{i})\otimes(\mathcal{L}_{\ell},0),$
which is a stable Higgs bundle with the exponent of nilpotency =p=p by the same argument as Lemma 4.2. By Lemma 2.3, there is no other gr-semistable filtration on $(\mathcal{V}_{\ell},\nabla_{\ell})$ such that the exponent of nilpotency of the associated graded Higgs bundle is $\leq p-1$. This completes the proof. ∎
@@
Let $C$ be a curve of genus $g\geq 2$. There exists a nilpotent semistable Higgs bundle $(\mathcal{E},\theta)$ of exponent $\leq 1$ and any big rank which is not strongly semistable.
@@
Set $(\mathcal{E},\theta)=(\mathcal{E}_{2(g-1)},\theta_{2(g-1)})$. Then $\mu(\mathcal{E})$ is an integer. Let $\mathcal{M}$ be a line bundle whose slope is the same as the slope of $(\mathcal{E},\theta)$. We know that $(\widetilde{\mathcal{E}},\widetilde{\theta})=(\mathcal{E},\theta)\oplus(\mathcal{M},0)$ is a nilpotent semistable Higgs bundle of exponent $\leq 1$. Since the inverse Cartier transform preserves direct sum,
@@
$0\subset F^{*}\mathcal{M}\subset F^{*}\mathcal{M}\oplus\mathcal{N}^{\bullet},$
@@
From now on, we use $K_{C}^{\frac{1}{2}}$ to denote a fixed line bundle such that its square is isomorphic to $K_{C}$.
@@
Theorem 4.3 ([LSYZ19, Theorem 2.1]).
@@
$(\mathrm{Gr}_{\mathrm{HN}}\circ C^{-1})(\mathcal{E}_{\operatorname{unif}},\theta_{\operatorname{unif}})\cong(\mathcal{E}_{\operatorname{unif}},\theta_{\operatorname{unif}})\otimes(\mathcal{L},0),$
where $\mathrm{HN}$ denotes for the Harder-Narasimhan filtration, and $\mathcal{L}$ is a torsion line bundle of order two.
Remark 4.1.
To be precise, in the statement of [LSYZ19, Theorem 2.1], the authors only mention the existence of such filtration, but it can be seen from the outline of proof that it is the Harder-Narasimhan filtration: In fact, the filtration is given as $\text{Fil}_{1}$ in the proof of Proposition $3.6$ in [LSYZ19].
Put $(\mathcal{V}_{\operatorname{unif}},\nabla_{\operatorname{unif}})=C^{-1}\circ\mathrm{Gr}_{\text{HN}}\circ C^{-1}(\mathcal{E}_{\operatorname{unif}},\theta_{\operatorname{unif}})$. Then by Theorem 4.3, it is a uniformizing flat bundle with nilpotent $p$-curvature. For the case $p=2$, the above theorem does not work. However, there is a direct way to construct a uniformizing flat bundle when $g(C)$ is odd: Since both $K_{C}\cong F^{*}K_{C}^{\frac{1}{2}}$ and $\mathcal{O}_{C}$ can be equipped with the canonical connection, we define $\mathcal{V}=K_{C}\oplus\mathcal{O}_{C}$ and equip it with a connection $\nabla$ defined by
(4.1) $\begin{pmatrix}\nabla_{\operatorname{can}}&\theta\\ 0&\nabla_{\operatorname{can}}\end{pmatrix}$,
@@
Clearly, $(\mathcal{E}_{1},\theta_{1})$ is a nilpotent stable Higgs bundle of exponent $\leq 1$. Next, let $\mathcal{E}_{2}=F_{*}K_{C}^{\frac{1-p}{2}}$ (note that $\mathcal{E}_{2}=F_{*}K_{C}^{-\frac{1}{2}}$ when $p=2$). By Theorem 4.1 and Riemann-Roch theorem, $\mathcal{E}_{2}$ is a stable bundle of rank $p$ and degree zero. Moreover, by Lemma 4.1, Lemma 4.2 and Lemma 2.2 the unique gr-semistable filtration on $F^{*}\mathcal{E}_{2}$ is given by its canonical filtration, and the associated graded Higgs bundle is
@@
where $\theta_{i}\colon K_{C}^{i}\to K_{C}^{i-1}\otimes K_{C}$ is an isomorphism for each $i=\frac{3-p}{2},\dots,\frac{p-1}{2}$. Third, we show that $(\mathcal{E}_{1},\theta_{1})\otimes(\mathcal{E}_{2},0)$ is a stable Higgs bundle. For $p=2$, if $(\mathcal{F},\theta_{\mathcal{F}})$ is a proper Higgs subbundle of $(\mathcal{E}_{1},\theta_{1})\otimes(\mathcal{E}_{2},0)$, its image in $(K_{C}\otimes\mathcal{E}_{2},0)$ would be of rank $\leq 1$ by the argument similar to Lemma 4.2 and then we can check that the slope of $\mathcal{F}$ is smaller than that of $\mathcal{E}_{1}\otimes\mathcal{E}_{2}$. For $p$ odd, we have explicitly
@@
which is a stable Higgs bundle: Indeed, if $(\mathcal{F},\theta_{\mathcal{F}})$ is a proper subbundle and if $r$ (resp. $s$) denotes the rank of the image (resp. kernel) of the projection from $(\mathcal{F},\theta_{\mathcal{F}})$ to $(\mathcal{E}_{2}\otimes K_{C}^{\frac{1}{2}},0)$, an argument similar to Lemma 4.2 implies the inequality $r\leq s$, and it implies that the slope of $\mathcal{F}$ is less than 0. Fourth, there is the following diagram of the Higgs-de Rham flow operator acting upon $(\mathcal{E}_{1},\theta_{1})\otimes(\mathcal{E}_{2},0)$:
$C^{-1}(\mathcal{E}_{1},\theta_{1})\otimes(F^{*}F_{*}K_{C}^{\frac{1-p}{2}},\nabla_{\operatorname{can}})$$(\mathcal{E}_{1},\theta_{1})\otimes(\mathcal{E}_{2},0)$$(\mathcal{E}^{\prime},\theta^{\prime})$$\scriptstyle{\mathrm{Gr}_{\mathcal{N}^{\bullet}}}$$\scriptstyle{C^{-1}}$
@@
By Lemma 4.4 below, $(\mathcal{E}^{\prime},\theta^{\prime})$ is polystable whose exponent of nilpotency is exactly $p$. Hence $\mathcal{N}^{\bullet}$ is a gr-polystable filtration. It follows from Lemma 2.3 that there is no gr-semistable filtration on $C^{-1}(\mathcal{E}_{1},\theta_{1})\otimes(F^{*}F_{*}K_{C}^{\frac{1-p}{2}},\nabla_{\operatorname{can}})$ such that its associated graded Higgs bundle has exponent of nilpotency $\leq p-1$. This shows that $(\mathcal{E}_{1},\theta_{1})\otimes(\mathcal{E}_{2},0)$ is not strongly semistable. Fifth, by the strong semistability theorem, both $(\mathcal{E}_{1},\theta_{1})$ and $(\mathcal{E}_{2},0)$ are strongly semistable. Summarizing the above altogether, we have constructed an example of strongly semistable Higgs bundles whose tensor product is semistable but not strongly semistable, as claimed.
@@
By the same argument as Lemma 4.2, we know that the $m$-th symmetric product of the uniformizing Higgs bundle
$\mathrm{Sym}^{m}(\mathcal{E}_{\operatorname{unif}},\theta_{\operatorname{unif}})=(K_{C}^{\frac{m}{2}}\oplus\dots\oplus K_{C}^{-\frac{m}{2}},\bigoplus_{i=1}^{m}\theta_{i})$
is stable.